OpenStudy (anonymous):

Tom's and Tim's ages add up to 21 years. After three years, Tom will be twice Tim’s age. How old are the two boys now? Tom: 12 years, Tim: 9 years Tom: 16 years, Tim: 5 years Tom: 14 years, Tim: 7 years Tom: 15 years, Tim: 6 years

OpenStudy (michele_laino):

I think that you problem can be modeled by the subsequent system: \[\left\{ \begin{gathered} to + ti = 21 \hfill \\ to + 3 = 2\left( {ti + 3} \right) \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

where to is Tom's age, whereas ti is Tim's age

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

we have to solve it. How many ways do you know for solving an algebraic system?

OpenStudy (anonymous):

for this problem i dont know

OpenStudy (michele_laino):

I rewrite the second equation as below: \[to + 3 = 2ti + 6\]

OpenStudy (michele_laino):

then I subtract 3, from both sides, so I get: \[to = 2ti + 3\]

OpenStudy (michele_laino):

now, we have to substitute that condition into the first equation, namely we have to substitute 2ti+3, in place of to into the first equation of our algebraic system, try please, what do you get?

OpenStudy (anonymous):

is it 5

OpenStudy (michele_laino):

what is 5?

OpenStudy (michele_laino):

hint: \[to + ti = \left( {2ti + 3} \right) + ti\]

OpenStudy (michele_laino):

so our first equation, can be rewritten as below: \[\left( {2ti + 3} \right) + ti = 21\]

OpenStudy (michele_laino):

please, solve it for "ti"

OpenStudy (michele_laino):

hint: next step: \[\begin{gathered} 2ti + 3 + ti = 21 \hfill \\ 3ti + 3 = 21 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

I subtract 3 from both sides, so I get: \[3ti = 18\]

OpenStudy (michele_laino):

what is "ti"?

OpenStudy (anonymous):

i really dont know

OpenStudy (michele_laino):

we have to divide both sides by 3, so we get: \[ti = \frac{{18}}{3} = ...?\]

OpenStudy (michele_laino):

therefore, the value of "ti" is: ti=...?

OpenStudy (anonymous):

wait is it 15 + 3 = 18

OpenStudy (michele_laino):

no, ti=6 so to=21-6=...?

OpenStudy (anonymous):

15

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

so what is the right option?

OpenStudy (anonymous):

d

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

thanks can you help with a couple more

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ok